OpenStudy (anonymous):

What is the solution to the equation? 4(3 + 7) = 4 • x + 4 • 7 A. x = 3 B. x = 4 C. x = 7 D. x = 10

jimthompson5910 (jim_thompson5910):

Notice how the 4 distributes (basically the outside 4 is multiplied by EVERY term inside) \[\Large 4(3+7) = 4 \cdot x + 4 \cdot 7\] \[\Large {\color{red}{4}}(3+7) = {\color{red}{4}} \cdot x + {\color{red}{4}} \cdot 7\]